2x+3y=21
-3x-6y=-24
: 2x+3y=21
-3x-6y=-24

Answer by Alan3354(1187)

2x+3y=21 eqn 1
-3x-6y=-24 eqn 2
----------
4x+6y=42 eqn 1 doubled
Add
x + 0y = 18
x = 18
Sub for x in eqn 1
2*18 + 3y = 21
36 + 3y = 21
3y = -15
y = -5